hello,,could u please answer my question??
The earning of three stalls, P,Q,R in a market are in the ratios ; P:Q = 3:4 and P:R = 5:6. Find the ratio of earnings of stall Q to stall R...

Answer by Alan3354(69443)

The earning of three stalls, P,Q,R in a market are in the ratios ; P:Q = 3:4 and P:R = 5:6. Find the ratio of earnings of stall Q to stall R
--------------
P/Q = 3/4
P = 3Q/4
-----
P/R = 5/6
P = 5R/6
----
3Q/4 = 5R/6
18Q = 20R
Q/R = 20/18 = 10/9